# ELUTQ: Efficient LUT-Aware Quantization for Deploying Large Language Models on Edge Devices

### v2

@@ Abstract
Weight quantization effectively reduces memory consumption and enable the deployment of Large Language Models on CPU-based edge devices, yet existing hardware-friendly methods often rely on uniform quantization, which suffers from poor weight-distribution fitting and high dequantization overhead under low-bit settings. In this paper, we propose ELUTQ, an efficient quantization framework featuring a novel quantization format termed Hierarchical Linear Quantization (HLQ). HLQ is designed to better capture the statistical characteristics of weights without increasing the computational cost of Bit-serial LUT-based GEMM operations, thereby eliminating dequantization overhead. HLQ is orthogonal to existing quantization algorithms. For the LLaMA3.1-8B model, when combined with the post-training quantization framework, HLQ enhances uniform quantization by achieving approximately 8% perplexity reduction at 3-bit precision and 85% perplexity reduction at 2-bit precision. When combined with efficient finetuning techniques, HLQ further improves model accuracy. A novel disk-offload technique is integrated into ELUTQ, enabling it to complete the quantization of LLaMA 3.1–70B using only 64 GB of CPU memory and 48 GB of VRAM, significantly reducing the hardware requirements for large-scale model quantization. To enable efficient deployment on edge devices, ELUTQ designs high-performance CPU kernels to support end-to-end inference. Under a 4-thread configuration with batch size = 1, our 2-bit quantized LLaMA2-7B model achieves a throughput of over 25 tokens per second on an Apple M2 chip. All the code is available at https://github.com/Nkniexin/ELUTQ.
Machine Learning, ICML
@@ 1 Introduction
Unlike high-performance GPU servers, on-device hardware typically operates under stringent resource constraints, characterized by limited memory capacity and computational power. These devices predominantly employ ARM or x86 CPUs, which offer restricted vectorization support and limited parallelism. To address these challenges, model quantization (Frantar et al., 2022; Xiao et al., 2023; Lin et al., 2024; Kim et al., 2023; Shang et al., 2023; Chen et al., 2024b) has emerged as a widely adopted compression technique, where high-precision weights are mapped to discrete integer values and stored using low-bit representations. This approach drastically reduces memory footprint while preserving model accuracy. Recent advances demonstrate that 8-bit weight quantization achieves near-lossless performance (Xiao et al., 2023; Yao et al., 2022). Furthermore, 4-bit quantization techniques (Frantar et al., 2022; Lin et al., 2024; Shao et al., 2023; Kim et al., 2023; Chen et al., 2024c) typically incur less than 3% accuracy degradation, with ongoing research further improving robustness. However, activation quantization remains more challenging due to the prevalence of outlier values. While some studies (Xiao et al., 2023; Liu et al., 2024) explore low-bit activation quantization, most practical implementations retain activations in FP16 precision (e.g., W4A16, W3A16) to ensure accuracy. Meanwhile, some research (Chen et al., 2024a; Shang et al., 2023; Huang et al., 2024) focuses on extreme low-bit weight quantization, pushing the boundaries of efficiency without sacrificing model quality.
Figure 1: Overview of ELUTQ.
@@
Quantization can be categorized as uniform (Frantar et al., 2022; Lin et al., 2024; Shao et al., 2023; Chen et al., 2024c) or non-uniform (Chee et al., 2023; Park et al., 2024; Kim et al., 2023; Xu et al., 2023; Zhao & Yuan, 2025) depending on whether the quantization intervals are equal. Uniform quantization divides the weight space into equal intervals, making it hardware-friendly and efficient for acceleration (Lin et al., 2024; Frantar et al., 2022). However, as noted in (Dettmers et al., 2023a; Kim et al., 2023), this approach poorly matches the bell-shaped distribution of typical weights, incurring substantial approximation error. Non-uniform quantization addresses this by adaptively allocating bins, via clustering (Kim et al., 2023; Zhao & Yuan, 2025) or codebooks (Chee et al., 2023) to better fit the weight distribution. While this improves representational efficiency, it often sacrifices hardware compatibility due to irregular memory access patterns. To address these limitations, We propose Hierarchical Linear Quantization, a non-uniform quantization format that reduces weight quantization error while maintaining hardware compatibility.
Many large model quantization methods aim to efficiently complete model quantization, where efficiency primarily refers to time and memory consumption, particularly video memory (VRAM), which is typically more constrained and expensive than CPU memory. Achieving quantization with low memory usage and short processing time is critical, as it enables running quantization algorithms even on consumer-grade GPUs such as the RTX 3090 or RTX 4090. Most post-training quantization approaches (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023; Huang et al., 2024; Shang et al., 2023; Zhao & Yuan, 2025; Kim et al., 2023) can complete quantization quickly and with modest memory requirements. These train-free methods operate without retraining and typically quantize the model layer by layer. For example, GPTQ quantizes one linear layer at a time. With sufficient CPU memory, even large models such as LLaMA3.1-70B can be quantized on a single A6000 GPU. In contrast, efficient finetuning–based methods (Shao et al., 2023; Dettmers et al., 2023b, a; Chee et al., 2023; Li et al., 2023; Xu et al., 2023; Chen et al., 2024b) require partial retraining, which increases time cost but often mitigates VRAM pressure by freezing most parameters, training only a small subset, or performing quantization in batches. This strategy allows quantization to be completed within limited video memory. However, while prior studies have primarily focused on reducing GPU memory consumption, the CPU memory overhead has been largely neglected. Existing quantization frameworks typically require loading the entire model into CPU memory during processing, which poses a critical bottleneck for large-scale models. For instance, quantizing a LLaMA 3.1–70B model requires at least 140 GB of CPU RAM, and the actual usage is often much higher, exceeding the capacity of most standard servers. In this paper, we integrate our proposed Hierarchical Linear Quantization with efficient quantization techniques to enhance model accuracy while maintaining low GPU memory usage, reduced CPU memory footprint, and short quantization time. For example, our framework can quantize the Llama 3.1-70B model using 2 RTX 4090 GPU (48GB VRAM) and only 64GB of CPU RAM, significantly reducing both GPU and CPU memory requirements.
@@
Although weight quantization reduces memory footprint, traditional methods require dequantization to higher precision (8-bit/FP16) for computation, introducing significant overhead that can paradoxically slow down inference at low bit-widths. Recent advances (Wei et al., 2025; Park et al., 2025b, 2022) address this by replacing standard GEMM with lookup table (LUT)-based operations that implement generalized FP-INT multiplication, eliminating dequantization while achieving both linear latency reduction with bit-width and improved energy efficiency. FLGLUT (Park et al., 2025b) accelerates these operations on GPUs and T-MAC (Wei et al., 2025) optimizes for CPUs via SIMD instructions. We enhance this paradigm through a pure C++ kernel redesign that specifically supports our novel Hierarchical Linear Quantization format, maintaining cross-platform compatibility while optimizing for our method’s unique requirements.
Figure  1 shows the ELUTQ design. Our contributions are as follows.
@@
  • We propose Hierarchical Linear Quantization (HLQ), a novel non-uniform quantization format that provides greater flexibility in weight representation compared to uniform quantization, while enabling efficient matrix computation through LUT-based GEMM.
  • We integrate HLQ into existing efficient quantization methods, including post-training quantization and efficient fine-tuning techniques, demonstrating that HLQ is orthogonal to most existing methods and can significantly enhance model performance under low-bit settings, without introducing noticeable memory or time overhead to the quantization pipeline.
  • We design an efficient CPU kernel tailored for the HLQ format, which enables high-performance matrix operations on edge devices.
  • We introduce ELUTQ, a unified quantization framework built upon HLQ. ELUTQ incorporates common quantization optimization techniques and includes a fully C++-implemented inference runtime, which supports end-to-end deployment of quantized models on edge devices.
@@ 2 Related works
2.1 Model Quantization
2.1.1 Uniform Quantization.
The uniform quantization formula is as follows:
Quantization:
${\mathrm{W}_{int}=\mathrm{clamp}}(\lfloor\frac{\mathrm{W}}{s}\rceil+z,0,2^{q}-1).$ (1)
Here, $\lfloor\cdot\rceil$ denotes the rounding operation, $q$ is the quantization bit width, $s$ is the scale factor, $z$ is the zero-point, $\mathrm{W}$ represents the original weights, and $\mathrm{W}_{int}$ denotes the quantized integer weights.
Dequantization:
$\mathrm{\hat{W}}=(\mathrm{W}_{int}\ -\ z)\cdot s,$ (2)
@@ 2.1 Model Quantization
where $\mathrm{\hat{W}}$ represents the dequantized weights. Uniform quantization is hardware-friendly, as dequantization can be implemented with simple multiplication. However, its representational capability is limited because it maps weights into a uniformly spaced range. Moreover, in modern computing systems, the smallest storage and computation unit is typically 8 bits. Therefore, for low-bit settings (e.g., 2-bit or 3-bit), the quantized integer weights $\mathrm{\hat{W}}$ must first be dequantized into 8-bit or 16-bit formats before computation, which often introduces non-negligible computational overhead.
2.1.2 Efficient Quantization.
Efficient quantization aims to complete model quantization with minimal time and memory consumption. It is generally categorized into two types: post-training quantization (PTQ) and efficient finetuning. PTQ requires only a small calibration dataset and can be completed within a relatively short time. Several research streams have advanced PTQ for Large Language Models. Some works, such as GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024), focus on weight-only quantization, demonstrating minimal performance degradation (e.g., less than 3% perplexity increase) with 4-bit uniform quantization. Other approaches (Dettmers et al., 2023b; Kim et al., 2023) differentiate weights by their significance, storing a subset of important weights in higher precision. Furthermore, studies including SqueezeLLM (Kim et al., 2023) and GANQ (Zhao & Yuan, 2025) observe that weights in LLMs often follow a bell-shaped distribution and subsequently propose non-uniform quantization methods based on k-means clustering. Beyond weight quantization, activation quantization has also been explored. Methods like SmoothQuant (Xiao et al., 2023) and LLM.int8() (Dettmers et al., 2022) have achieved notable results in the challenging W8A8 weight-activation joint quantization setting. Compared with post-training quantization, efficient finetuning typically requires a little more data and time to search and adjust model parameters. OmniQuant (Shao et al., 2023) searches quantization parameters block by block and is the first to achieve promising results under 2-bit settings. PB-LLM (Shang et al., 2023) employs a feature segmentation strategy to achieve competitive performance under 2-bit weight quantization. Meanwhile, DB-LLM (Chen et al., 2024b) decomposes 2-bit quantized weights into two independent sets of binary weights and further utilizes distillation to enhance model performance, albeit at the cost of introducing substantial fine-tuning overhead. BiLLM (Huang et al., 2024) pushes the boundary further by leveraging weight distribution characteristics to achieve an average bit-width of approximately 1.11 bits.
Figure 3: Detail pipeline of Bit-serial Lut-based GEMM.
2.2 LUT-Based GEMM
There are two primary computational paradigms for LUT-based GEMM, as illustrated in the Figure 2.
@@ 2.2 LUT-Based GEMM
Figure 2(c) illustrates another GEMM paradigm termed Bit-serial LUT-based GEMM. Unlike the scheme in Figure 2(b), this method utilizes the LUT to store all possible dot products between an activation vector and single-bit weights. The detail computational procedure is shown in Figure 3. First, a $q$-bit quantized weight matrix $\mathrm{W}_{int}$ is decomposed into $q$ single-bit matrices $\{W_{0},W_{1},...,W_{q-1}\}$ offline, where each element is either 0 or 1, representing the respective bit planes of the original weights. For example, for integer values (9, 7, 6, 3) with binary representations (1001, 0111, 0110, 0011), the matrix for the lowest bit is (1, 1, 0, 1), and the matrix for the highest bit is (1, 0, 0, 0). This decomposition is performed offline, incurring no runtime overhead. During inference, for an activation vector of the group size $g$, the system precomputes the dot products between this activation vector and all $2^{g}$ possible combinations of single-bit weights, storing the results in the LUT. Thus, the original matrix computation requiring high-precision multiply-accumulate operations is simplified into highly efficient table lookups followed by summation. This paradigm has been demonstrated to offer high computational efficiency and energy efficiency (Park et al., 2025b; Wei et al., 2025). For example, FIGLUT (Park et al., 2025b) optimized the table structure for GPU architectures to avoid bank conflicts, while T-MAC leveraged CPU vectorized lookup instructions (AVX2/NEON) to enable efficient LUT operations on CPUs.
The HLQ method is built upon the Bit-serial LUT-based GEMM computational paradigm. In contrast to prior works like FIGLUT and T-MAC, which primarily focused on designing efficient computational kernels for this paradigm, our work emphasizes optimization at the quantization algorithm level. The objective is to enhance the accuracy of low-bit quantized models, thereby achieving a synergistic improvement in both accuracy and efficiency within this highly efficient computational paradigm.
2.3 Inference Framework
Many frameworks aim to enable efficient inference of quantized models across a wide range of hardware platforms. For GPUs, vLLM (Kwon et al., 2023) employs the PageAttention technique to optimize key-value (KV) memory management and supports quantization formats such as GPTQ and AWQ. SGlang (Zheng et al., 2024) reduces response latency through shared prefix requests and efficient caching strategies. In addition, frameworks like TensorRT-LLM (NVIDIA, 2023) and MLC-LLM (Chen, 2023) have also been developed for GPU-optimized inference. For CPUs and other edge processors, llama.cpp (Gerganov, 2023) is a lightweight framework implemented entirely in C++. Although its inference speed is slower compared to GPU-accelerated solutions and thus not suitable for large-scale online services, it is well-suited for edge computing, IoT, and low-throughput scenarios, providing a practical solution for basic inference in GPU-free environments.
@@ 3 Motivation
3.1 Bell-Shaped Distribution of Weights
As noted in prior studies (Kim et al., 2023; Dettmers et al., 2023a; Zhao & Yuan, 2025), weights often exhibit a bell-shaped distribution. Uniform quantization, which maps weights to a uniformly spaced grid, struggles to accurately capture such distributions, especially under low-bit settings. Figure 4 shows that the weight distribution of an out_channel in the up_projection layer of the LLaMA3.1-8B model closely follows a Gaussian pattern, with values heavily concentrated near zero. This characteristic, commonly observed across various models, highlights the inherent limitation of uniform quantization in effectively representing natural weight distributions. Consequently, non-uniform quantization strategies are often necessary to achieve higher fidelity in low-bit compression.
Figure 4: (Top) The weight distribution of one output channel in a up_proj of LLaMA3.1-8B. (Bottom) MSE loss for weight comparsion between Uniform and HLQ.
3.2 Optimization for Bit-Serial LUT-Based GEMM
For ordinary LUT-based GEMM algorithms shown in figure 2(b), previous efforts such as SqueezeLLM (Kim et al., 2023) and GANQ (Zhao & Yuan, 2025) have achieved significant improvements in algorithmic performance for 3-bit and 4-bit quantization, along with the development of highly efficient CUDA kernels for accelerated inference. However, these implementations still lack support for CPU devices. Regarding Bit-serial LUT-Based GEMM, existing works including T-MAC (Wei et al., 2025), LUT-GEMM (Park et al., 2022), and FIGULT (Park et al., 2025b) have primarily focused on optimizing kernel design to support efficient uniform quantization schemes, yet they do not address algorithmic enhancements for improving quantization accuracy. Contemporary work like AnyBCQ (Park et al., 2025a) has made certain explorations, focusing on designing multi-precision models, with its 2-bit models achieving limited performance and primarily targeting model inference on GPUs. This limitations restricts the broader application of such kernels in cpu-based edge devices. To bridge this gap, our work proposes using HLQ to boost quantization accuracy at the algorithm level, while leveraging Bit-Serial LUT-Based GEMM to enable efficient inference on CPU devices.
3.3 Why Choose Bit-Serial LUT-Based GEMM for Edge Devices
Edge devices are predominantly based on CPU architectures, which offer significantly lower programmibility compared to GPU architectures. This limitation prevents developers from customizing lookup tables to implement LUT-based GEMM shown in figure 2(b). However, CPUs with x86_64 or ARM architectures provide a set of vectorized table-lookup instructions, such as _mm256_shuffle_epi8 and vqtbl1q_u8, which naturally align with the computational paradigm of Bit-serial LUT-based GEMM. Prior work such as T-MAC (Wei et al., 2025) has demonstrated that these instructions can be leveraged to enable efficient execution of Bit-serial LUT-based GEMM on edge devices. This observation motivates our work to bridge the algorithm-kernel co-design gap by combining high-accuracy non-uniform quantization at the algorithm level with a Bit-serial LUT-based inference framework tailored for common CPU instruction sets.

### v3

# ELUTQ: Optimizing Quantization Accuracy under LUT-Based Computation for Edge LLMs

@@ Abstract
Weight quantization effectively reduces memory consumption and enable the deployment of Large Language Models on edge devices, yet existing hardware-friendly methods often rely on uniform quantization, which suffers from poor weight-distribution fitting and high dequantization overhead under low-bit settings. In this paper, we propose ELUTQ, an efficient quantization framework featuring a novel quantization format termed Hierarchical Linear Quantization (HLQ). HLQ is designed to better capture the statistical characteristics of weights and eliminate dequantization overhead using Bit-serial LUT-based GEMM operations. HLQ significantly improves model accuracy under low-bit settings and achieves performance comparable to QAT methods without any retraining of the weights. Moreover, an optimized quantization pipeline is integrated into ELUTQ, enabling it to complete the quantization of LLaMA 3.1–70B using only 64 GB of CPU memory and 48 GB of VRAM, reducing the hardware requirements for large-scale model quantization. To enable efficient deployment on edge devices, ELUTQ designs high-performance kernels to support end-to-end inference. Our 2-bit LLaMA3.1-8B achieves 1.5× speedup over AWQ on RTX 3090. Code is available at https://github.com/Nkniexin/ELUTQ.
footnotetext: * denotes the corresponding author.
Figure 1: (a) ELUTQ outperforms uniform quantization methods but remains slightly behind codebook-based non-uniform approaches. (b) ELUTQ achieves high efficiency in both quantization time and memory usage on RTX4090s when quantizing LLaMA3.1-70B. (c) ELUTQ leverages Bit-serial LUT-based GEMM operator to significantly accelerate low-bit inference on RTX3090.
@@ 1 Introduction
On-device hardware typically operates under stringent resource constraints, characterized by limited memory capacity and computational power. To address these challenges, model quantization Frantar et al. (2022); Xiao et al. (2023); Lin et al. (2024); Kim et al. (2023); Shang et al. (2023); Chen et al. (2024a) has emerged as a widely adopted compression technique, where high-precision weights are mapped to discrete integer values and stored using low-bit representations. This approach drastically reduces memory footprint while preserving model accuracy. Recent advances demonstrate that 8-bit weight quantization achieves near-lossless performance Xiao et al. (2023); Yao et al. (2022). Furthermore, 4-bit quantization techniques Frantar et al. (2022); Lin et al. (2024); Shao et al. (2023); Kim et al. (2023); Chen et al. (2024b) typically incur less than 3% accuracy degradation, with ongoing research further improving robustness.
@@
Quantization can be categorized as uniform Frantar et al. (2022); Lin et al. (2024); Shao et al. (2023); Chen et al. (2024b) or non-uniform Chee et al. (2023); Park et al. (2024); Kim et al. (2023); Xu et al. (2023); Zhao and Yuan (2025) depending on whether the quantization intervals are equal. Uniform quantization divides the weight space into equal intervals, making it hardware-friendly and efficient for acceleration Lin et al. (2024); Frantar et al. (2022). However, as noted in Dettmers et al. (2023a); Kim et al. (2023), this approach poorly matches the bell-shaped distribution of typical weights, incurring substantial approximation error. Non-uniform quantization addresses this by adaptively allocating bins, via clustering Kim et al. (2023); Zhao and Yuan (2025) or codebooks Chee et al. (2023) to better fit the weight distribution. Although this reduces approximation error, it often sacrifices hardware compatibility due to irregular memory access patterns. To address these limitations, we propose Hierarchical Linear Quantization, a non-uniform quantization scheme that achieves a trade-off between accuracy and hardware efficiency.
Many large model quantization methods aim to efficiently complete model quantization, where efficiency primarily refers to time and memory consumption. Prior studies have primarily focused on reducing GPU memory consumption, the CPU memory overhead has been largely neglected. For instance, Quip# Tseng et al. (2024) requires at least 1,000 GB of CPU memory and over 200 hours to quantize LLaMA-3.1-70B, exceeding the capacity of most standard servers. In this paper, for HLQ, we redesign the quantization pipeline to perform quantization under strict resource constraints, maintaining low GPU memory usage, a small CPU memory footprint, and short quantization time. For example, our framework can quantize the Llama 3.1-70B model using only 48GB VRAM and 64GB of CPU RAM in 40 hours.
@@
Although weight quantization reduces memory footprint, traditional methods require dequantizing low-bit weights to higher precision (e.g., INT8 or FP16) before computation. This dequantization incurs substantial unpacking overhead, which can paradoxically slow down inference at very low bit-widths. Recent advances Wei et al. (2025) address this by replacing standard GEMM with lookup table (LUT)-based operations that implement generalized FP-INT multiplication, eliminating dequantization while achieving both linear latency reduction with bit-width and improved energy efficiency. FLGLUT Park et al. (2025b) accelerates these operations on GPUs and T-MAC Wei et al. (2025) optimizes for CPUs via SIMD instructions. We enhance this paradigm through a pure C++ kernel redesign that specifically supports our novel Hierarchical Linear Quantization format, maintaining cross-platform compatibility.
Our contributions are as follows.
@@
  • Hierarchical Linear Quantization. A novel non-uniform quantization format that provides greater flexibility in weight representation compared to uniform quantization, while preserving hardware-friendly computation.
  • Efficient Quantization Pipeline for HLQ. We design an efficient quantization pipeline tailored to the characteristics of the HLQ, enabling model quantization with significantly reduced memory usage and runtime.
  • High-Performance LUT-Based GEMM Kernels. We design high-performance LUT–based GEMM kernels for HLQ on both CPUs and GPUs, enabling end-to-end inference across a wide range of edge devices.
@@ 2 Related works
Uniform and Non-Uniform Quantization. Uniform quantization Frantar et al. (2022); Lin et al. (2024); Shao et al. (2023) is hardware-friendly, as dequantization can be implemented with simple multiplication. However, its representational capability is limited because it maps weights into a uniformly spaced range. Codebook- and clustering-based non-uniform quantization methods Tseng et al. (2024); Egiazarian et al. (2024); Kim et al. (2023); Zhao and Yuan (2025) have demonstrated strong effectiveness in modeling complex weight distributions, often yielding lower representation error than uniform quantizers. Therefore, at low bit-widths, non-uniform quantization can achieve better performance than uniform quantization. However, in modern computing systems, the smallest storage and computation unit is typically 8 bits. Therefore, for low-bit settings (e.g., 2-bit or 3-bit), the quantized integer weights must first be dequantized into 8-bit or 16-bit formats before computation, which often introduces non-negligible unpacking overhead.
Efficient Quantization. Efficient quantization aims to complete model quantization with minimal time and memory consumption. For example, GPTQ Frantar et al. (2022) quantizes one linear layer at a time. With sufficient CPU memory, even large models such as LLaMA3.1-70B can be quantized on a single A6000 GPU. In contrast, efficient finetuning–based methods Shao et al. (2023); Dettmers et al. (2023b, a); Chee et al. (2023); Li et al. (2023); Xu et al. (2023); Chen et al. (2024a) require partial retraining, which increases time cost but often mitigates VRAM pressure by freezing most parameters, training only a small subset, or performing quantization in batches. These strategies make quantization to be completed within limited VRAM.
Figure 2: Illustration of two GEMM computation paradigms for weight-only quantized matrix multiplication. We provide a detailed explanations in Appendix B.
@@
Bit-Serial LUT-Based GEMM. Figure 2(b) illustrates a GEMM paradigm termed Bit-serial LUT-based GEMM. Unlike the dequant-based scheme in figure 2(a), this method utilizes the LUT to store all possible dot products between an activation vector and single-bit weights. Then, a $q$-bit quantized weight matrix $\mathrm{W}_{int}$ is decomposed into $q$ single-bit matrices $\{\mathrm{W}_{0},\mathrm{W}_{1},...,\mathrm{W}_{q-1}\}$ offline, where each element is either 0 or 1, representing the respective bit planes of the original weights. Thus, the original matrix computation requiring high-precision multiply-accumulate operations is simplified into highly efficient table lookups followed by summation (see detailed computational procedure in Appendix B). This paradigm has been demonstrated to offer high computational efficiency and energy efficiency Park et al. (2025b); Wei et al. (2025). The HLQ method is built upon the Bit-serial LUT-based GEMM computational paradigm.
@@ 3 Motivation
3.0.1 Bell-Shaped Distribution of Weights
As noted in prior studies Dettmers et al. (2023a); Kim et al. (2023), weights often exhibit a bell-shaped distribution. Uniform quantization, which maps weights to a uniformly spaced grid, struggles to accurately capture such distributions, especially under low-bit settings. Figure 3 shows that the weight distribution of an out_channel in the up_projection layer of the LLaMA3.1-8B model closely follows a Gaussian pattern, with values heavily concentrated near zero. This characteristic, commonly observed across various models, highlights the inherent limitation of uniform quantization in effectively representing natural weight distributions.
3.0.2 Hardware Inefficiencies in Codebook and Clustering Quantization
Methods based on codebooks (e.g., Quip#, AQLM) or clustering (e.g., SqueezeLLM, GANQ) generally lack hardware friendliness due to irregular memory access patterns. While these approaches capture weight distributions more accurately than uniform quantization, their reliance on indirection tables or cluster lookups introduces significant computational overhead. As shown in the figure  8, this overhead leads to even worse inference performance than FP16, resulting in substantial degradation in inference speed.
3.0.3 Optimization for Bit-Serial LUT-Based GEMM
Regarding Bit-serial LUT-Based GEMM, existing works including T-MAC Wei et al. (2025), LUT-GEMM Park et al. (2022), and FIGULT Park et al. (2025b) have primarily focused on optimizing kernel design, yet they do not address algorithmic enhancements for improving quantization accuracy. Contemporary work like AnyBCQ Park et al. (2025a) has made certain explorations, focusing on designing multi-precision models, with its 2-bit models achieving limited performance and primarily targeting model inference on GPUs. This limitations restricts the application of such kernels in broader scenarios. To bridge this gap, our work proposes HLQ to boost quantization accuracy at the algorithm level, while leveraging Bit-Serial LUT-Based GEMM to enable efficient inference on edge devices, including CPUs and GPUs.
Figure 3: (Top) The weight distribution of one output channel in a up_proj of LLaMA3.1-8B. (Bottom) Mean squared error between quantized weights and original weights for Uniform quantization and HLQ.